OpenStudy (anonymous):

a tearcher spent $64 buying two types of games. one type cost $12 each and the other $14 each. how many of the $12 games did she bye?

OpenStudy (anonymous):

Let us say she bought T of the $12 one and F of the $14 one. Equation says 12T + 14F = 64.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Need more information to solve this problem because you can't solve two variables with a single equation.

OpenStudy (anonymous):

na u can

OpenStudy (anonymous):

with the above equation only one logical possiblities is possible!!

OpenStudy (anonymous):

Oh sure....but, that is the "second equation". You can guess about it. But, it is the second equation.

OpenStudy (anonymous):

try putting some values less than 5

OpenStudy (anonymous):

Trying random guesses is not same as solving. In this case, you need a second equation.

OpenStudy (anonymous):

@GT see t<5 f<5 itz obvious as 12*5=60 14*5=70

OpenStudy (anonymous):

fr such ques. u 've 2 substitute logical possiblities

OpenStudy (anonymous):

It appears to me that you are trying to convince me as opposed to teaching/answering the original poster.

OpenStudy (anonymous):

see such questions ...should be answered logical way ...*no offence!!

OpenStudy (anonymous):

and in fact it is easy to find!!

OpenStudy (anonymous):

I understand. And, you are a smart dude.

OpenStudy (anonymous):

lol ....thanks ... and was dat out of sarcasm ... ;) * jk!!

OpenStudy (anonymous):

No. I know you are. The point is to explain and teach the person who actually posted the question. Or, in worst case offer an answer to that person.